A hot liquid at 80 ̊ C is added to 600 g of the same liquid originally at 10 ̊ C. When the
mixture reaches 30 ̊ C, what will be the total mass of the liquid?
1
Expert's answer
2020-07-02T17:52:19-0400

"T_1 = 80^\circ C"

"T_2 = 10^\circ C"

"T_0 = 30^\circ C"

"M_2 = 600 g"

"M_1 +M_2 = ?"


"\u0441_1M_1(T_0-T_1) +c_2M_2(T_0-T_2) = 0" "c_1=c_2" (the same liquid)


Then "M_2(T_0 - T_2) = M_1(T_1 - T_0)" "\Rightarrow M_1 = \dfrac{M_2(T_0 - T_2)}{T_1 - T_0} = 600g \cdot \dfrac{30 - 10}{80 - 30} = 240g"

Finally, "M_1 +M_2 = 240g + 600g = 840g"